OpenStudy (anonymous):

Calculate the discriminant to determine the number of real roots. y = x2 + 3x + 9 How many real roots does the equation have? A. two real roots B. no real roots C. one real root D. no solution to the equation

OpenStudy (anonymous):

@LynFran

OpenStudy (amoodarya):

\[\sqrt{b^2-4ac}=\sqrt{3^2-4*9*1}\]

OpenStudy (anonymous):

so A there are two real roots?

OpenStudy (solomonzelman):

Ok, I will tell you this. The discriminant for any quadratic in a form of: \(\color{black}{ \displaystyle y=a{\rm x}^2+b{\rm x}+c }\) is: \(\large\color{black}{ \displaystyle {\rm D}=\sqrt{b^2-4ac} }\)

OpenStudy (solomonzelman):

So, if the discriminant is: an integer - then you can factor the quadratic a real number (not 0, and not integer) - then you just have 2 real roots. an imaginary number (that results from a negative in a square root) - then no real roots, rather, only 2 imaginary roots.

OpenStudy (anonymous):

for people that see this in the future its B